[Brahms-l] Shift report 20060620 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Jun 20 2006 - 08:12:40 EDT
Shifter leader was ps
Shifters were nt


<pre>
We run from #16274 (previous shift) through #16281 (next shift).
The condition is
FS 2.3 deg 1723 A(D1,D2,D3,D4) + MRS 90 deg 350 A

Run #16274 (from previous shift)

The beam was dumpped at around 01:00 am.
#16275 as TPC pedestal run.
#16276 as Non-TPC pedestal run.

Run from #16277 through #16280.
Run #16281 (to next shift)
There is no detector problems so far.
</pre>
_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jun 20 08:13:07 2006

This archive was generated by hypermail 2.1.8 : Tue Jun 20 2006 - 08:13:22 EDT